# Boltzmann Generators for condensed matter via Riemannian Flow Matching

### v1

@@ A.2 Systems
A.2.1 3N-dim Gaussian prior: Einstein-Crystal
@@
where $\Lambda$ is the de Broglie wavelength, which is commonly set to a characteristic length scale of the interaction potential in model systems. The second term accounts for the constraint of fixing the center of mass (Frenkel and Ladd (1984)).
By computing the free energy difference between the Einstein Crystal and any arbitrary system, one can obtain the absolute free energy of the target system, which allows to rank different phases of a given system by stability. We follow Wirnsberger et al. (2022) by setting the width of the Gaussians, which is related to $\Lambda_{E}$, similar to the width of the displacements of the atoms around their equilibrium positions. For us, this ensures that the vector field always acts on a similar environment for each time $t$. The widths for the different systems are collected in Tab. 3.
@@ A.2.2 Monoatomic-Water
The monoatomic water is a special case of the Stillinger-Weber (SW) potential Stillinger and Weber (1985). It incorporates two-body terms ($\phi_{2}$) as well as three-body ($\phi_{3}$) interaction terms, where the latter promote tetrahedral coordination environments. The total potential energy is expressed as:
@@
Simulations were performed with the same settings as in Wirnsberger et al. (2022) and Schebek et al. (2025b) in the NVTNVT ensemble using the openMM package (Eastman et al. (2023)). The mW ice was modeled at $T=200$ K at a density of $\rho=1.004$ g cm-3.
@@ A.3 Second-Order Cumulant Expansion for Hutchinson’s Estimator
Continuous normalizing flows (CNFs) transform samples $\mathbf{x}\sim p(\mathbf{x})$ according to
$f^{t}_{\theta}(\mathbf{x})=\mathbf{x}+\int_{0}^{t}\,v_{\theta}\!\left(t^{\prime},\,f^{t^{\prime}}_{\theta}(\mathbf{x})\right)dt^{\prime}$ (16)
where $v_{\theta}$ is a time-dependent vector field. The generated samples $f^{1}_{\theta}(\mathbf{x})$ are distributed according to the pushforward distribution $\tilde{p}$. The associated change in log-density is given by the instantaneous change of variables formula:
$\Delta\log p:=\log\tilde{p}(f^{1}_{\theta}(\mathbf{x})))-\log p(\mathbf{x})=-\int_{0}^{1}\operatorname{Tr}(\nabla v_{\theta}(t,f_{\theta}^{t}(\mathbf{x})))\,dt,$ (17)
where $\nabla v_{\theta}$ denotes the Jacobian of the vector field, the trace of which is the divergence. The computational cost of the CNF vector field evaluation scales linearly with system size $\mathcal{O}(N)$, due to the local cutoff of the GNN. Calculating the exact divergence requires the trace of the full Jacobian, which scales linearly with the input dimension $D\propto N$, resulting in an overall $\mathcal{O}(N^{2})$ complexity for both computation and memory, rendering the computation of the exact trace infeasible for large systems. To scale to large systems, we employ Hutchinson’s trace estimator. At each integration step $t$, we approximate the trace using $K$ stochastic noise vectors $\{\bm{\epsilon}_{k}\}_{k=1}^{K}$ with Rademacher entries, i.e., $\epsilon_{i}\in\{-1,+1\}$ i.i.d. with equal probability 1/2:
@@
$\widehat{\log w(\mathbf{x})}=\beta_{p}U_{p}(\mathbf{x})-\beta_{\mu}U_{\mu}(f^{1}_{\theta}(\mathbf{x}))+\widehat{\Delta\log p}$. (19)
However, the computation of thermodynamic observables, such as the free energy difference, depend on the exponentiated log weights. By applying Jensen’s inequality to the expectation over the Hutchinson probes $\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}$, we see that:
$\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\hat{w}(\mathbf{x})]=C(\mathbf{x})\cdot\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\exp(\widehat{\Delta\log p})]\geq C(\mathbf{x})\cdot\exp(\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\widehat{\Delta\log p}])=w(\mathbf{x}),$ (20)
where $C(\mathbf{x})$ contains the deterministic potential energy terms. This inequality implies that for any given sample $\mathbf{x}$, the stochastic weight is systematically overestimated on average. This bias propagates to the free energy estimator, which involves a second expectation over the samples $\mathbf{x}$ drawn from the prior $p(\mathbf{x})$. The estimator for the exponential of the free energy difference behaves as follows:
$\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\hat{w}(\mathbf{x})]\right]\geq\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}[w(\mathbf{x})]=e^{-\beta\Delta F}.$ (21)
@@
$\beta\Delta\hat{F}\coloneq-\log\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x}),\bm{\epsilon}\sim p(\bm{\epsilon})}[\hat{w}(\mathbf{x})]\leq-\log\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}[w(\mathbf{x})]=\beta\Delta F.$ (22)
To address this, we treat the stochastic error from the trace estimator as a fluctuating work term and apply the second-order cumulant expansion, a method rooted in thermodynamic perturbation theory Zwanzig (1954). The expected value of the likelihood estimate can be expressed via the cumulant generating function expansion:
@@
$=\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\widehat{\Delta\log p}]+\frac{1}{2}\bigg(\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[(\widehat{\Delta\log p})^{2}]-\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\widehat{\Delta\log p}]^{2}\bigg)+\frac{1}{6}$... (24)
We truncate this series after the second order. This approximation is justified by the Central Limit Theorem: because the total accumulated noise arises from the sum of independent stochastic evaluations over many integration steps, its distribution converges rapidly to a Gaussian, even if the individual noise terms are non-Gaussian. Since all cumulants $\kappa_{n}$ with $n>2$ vanish for a Gaussian distribution, we correct the bias in the log-weights using the accumulated variance $\hat{\sigma}^{2}$ of the trace estimator along the trajectory. Since each sample $\mathbf{x}$ experiences a unique realization of the stochastic noise, both the log-density estimate $\widehat{\Delta\log p}$ and the variance correction $\hat{\sigma}^{2}$ are sample-specific:
$\widehat{\log w_{\text{corr}}}(\mathbf{x})\coloneq\widehat{\log w}(\mathbf{x})-\frac{1}{2}\hat{\sigma}^{2}(\mathbf{x}).$ (25)
@@
Finally, the stochasticity of the trace estimator directly impacts sampling efficiency. Since the importance weights inherit the noise from the $\Delta\log p$ estimate, the ESS inevitably degrades when fewer probes are used. Figure 3 demonstrates that with a sufficient number of probes, the ESS converges to the baseline value obtained from exact divergence calculations. It is crucial to note, however, that the validity of our correction relies on an accurate estimate of the noise variance $\hat{\sigma}^{2}$. Using too few probes may yield an unreliable variance estimate, which can distort the correction term itself and result in residual bias.
@@ A.5 Implementation details
The work was implemented in python, using pytorch Paszke et al. (2019), lightning Falcon and The PyTorch Lightning team (2019), pytorch-geometric Fey et al. (2025), and POT Flamary et al. (2021). To avoid waiting for mini-batch optimal transport reordering during training, it was reordered on the CPU simultaneously with training on the GPU. The code is available under https://github.com/emil-ho/bg-cnf-fm-condensed-matter.

### v2

@@ A.2 Systems
A.2.1 3NN-dim Gaussian prior: Einstein-Crystal
@@
where $\Lambda$ is the de Broglie wavelength, which is commonly set to a characteristic length scale of the interaction potential in model systems. The second term accounts for the constraint of fixing the center of mass (Frenkel and Ladd, 1984).
By computing the free energy difference between the Einstein Crystal and any arbitrary system, one can obtain the absolute free energy of the target system, which allows ranking different phases of a given system by stability. We follow Wirnsberger et al. (2022) by setting the width of the Gaussians, which is related to $\Lambda_{E}$, similar to the width of the displacements of the atoms around their equilibrium positions, with the exact widths provided in Tab. 3. For us, this ensures that the vector field always acts on a similar environment for each time $t$. The widths for the different systems are collected in Tab. 3.
@@ A.2.2 Monoatomic-Water
The monoatomic water is a special case of the Stillinger-Weber (SW) potential (Stillinger and Weber, 1985). It incorporates two-body terms ($\phi_{2}$) as well as three-body ($\phi_{3}$) interaction terms, where the latter promote tetrahedral coordination environments. The total potential energy is expressed as:
@@
Simulations were performed with the same settings as in Wirnsberger et al. (2022) and Schebek et al. (2025b) in the $NVT$ ensemble using the openMM package (Eastman et al., 2023). The mW ice was modeled at $T=200$ K at a density of $\rho=1.004$ g cm-3.
@@ A.3 Second-Order Cumulant Expansion for Hutchinson’s Estimator
To scale to large systems, we circumvent the effective $\mathcal{O}(N^{2})$ complexity of the exact divergence calculation by employing Hutchinson’s trace estimator. At each integration step $t$, we approximate the trace using $K$ stochastic noise vectors $\{\bm{\epsilon}_{k}\}_{k=1}^{K}$ with Rademacher entries, i.e., $\epsilon_{k,i}\in\{-1,+1\}$ i.i.d. with equal probability 1/2:
@@
$\widehat{\log w}(\mathbf{x}_{0})=\beta_{p}U_{p}(\mathbf{x}_{0})-\beta_{\mu}U_{\mu}(f_{\theta}(\mathbf{x}_{0}))-\widehat{\Delta\log p}$. (20)
However, the computation of thermodynamic observables, such as the free energy difference, depends on the exponentiated log weights. By applying Jensen’s inequality to the expectation over the Hutchinson probes $\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}$, we see that:
$\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\hat{w}(\mathbf{x}_{0})]=C(\mathbf{x}_{0})\cdot\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\exp(\widehat{\Delta\log p})]\geq C(\mathbf{x}_{0})\cdot\exp(\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\widehat{\Delta\log p}])=w(\mathbf{x}_{0}),$ (21)
where $C(\mathbf{x}_{0})$ contains the deterministic potential energy terms. This inequality implies that for any given sample $\mathbf{x}_{0}$, the stochastic weight is systematically overestimated on average. This bias propagates to the free energy estimator, which involves a second expectation over the samples $\mathbf{x}_{0}$ drawn from the prior $p(\mathbf{x}_{0})$. The estimator for the exponential of the free energy difference behaves as follows:
$\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0})}\left[\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\hat{w}(\mathbf{x}_{0})]\right]\geq\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0})}[w(\mathbf{x}_{0})]=e^{-\beta\Delta F}.$ (22)
@@
$\beta\Delta\hat{F}\coloneq-\log\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0}),\bm{\epsilon}\sim p(\bm{\epsilon})}[\hat{w}(\mathbf{x}_{0})]\leq-\log\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0})}[w(\mathbf{x}_{0})]=\beta\Delta F.$ (23)
To address this, we treat the stochastic error from the trace estimator as a fluctuating work term and apply the second-order cumulant expansion, a method rooted in thermodynamic perturbation theory (Zwanzig, 1954). The expected value of the likelihood estimate can be expressed via the cumulant generating function expansion:
@@
$=\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\widehat{\Delta\log p}]+\frac{1}{2}\bigg(\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[(\widehat{\Delta\log p})^{2}]-\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[\widehat{\Delta\log p}]^{2}\bigg)+$... (25)
We truncate this series after the second order. This approximation is justified by the Central Limit Theorem: because the total accumulated noise arises from the sum of independent stochastic evaluations over many integration steps, its distribution converges rapidly to a Gaussian, even if the individual noise terms are non-Gaussian. Since all cumulants $\kappa_{n}$ with $n>2$ vanish for a Gaussian distribution, we correct the bias in the log-weights using the accumulated variance $\hat{\sigma}^{2}$ of the trace estimator along the trajectory. Since each sample $\mathbf{x}_{0}$ experiences a unique realization of the stochastic noise, both the log-density estimate $\widehat{\Delta\log p}$ and the variance correction $\hat{\sigma}^{2}$ are sample-specific:
$\widehat{\log w_{\text{corr}}}(\mathbf{x}_{0})\coloneq\widehat{\log w}(\mathbf{x}_{0})-\frac{1}{2}\hat{\sigma}^{2}(\mathbf{x}_{0}).$ (26)
@@
Finally, the stochasticity of the trace estimator directly impacts sampling efficiency. Since the importance weights inherit the noise from the $\Delta\log p$ estimate, the ESS inevitably degrades when fewer probes are used. Figure 3 demonstrates that with a sufficient number of probes, the ESS converges to the baseline value obtained from exact divergence calculations. It is crucial to note, however, that the validity of our correction relies on an accurate estimate of the noise variance $\hat{\sigma}^{2}$. Using too few probes may yield an unreliable variance estimate, which can distort the correction term itself and result in residual bias. We note that the ESS is negligibly affected by the bias-correction.
@@ A.5 Implementation details
The work was implemented in python, using pytorch (Paszke et al., 2019), lightning (Falcon and The PyTorch Lightning team, 2019), pytorch-geometric (Fey et al., 2025), and POT (Flamary et al., 2021). To avoid waiting for mini-batch optimal transport reordering during training, it was reordered on the CPU simultaneously with training on the GPU. The code is available under https://github.com/emil-ho/bg-cnf-fm-condensed-matter.